You are given the following data for the year 2016 of XYZ Co. Limited

Variable cost              600,000

Fixed cost                  300000

Net profit                   100,000

1000,000

Required, Find out:

a)    Break-even point

b)    P/V ratio

c)    Margin of safety

d) Draw a break even chart showing contribution 


1
Expert's answer
2022-03-22T18:35:01-0400

Contribution: Fixed cost + Profit

= 300000+100000

=400000


Sales: Contribution + Variable cost

=400000+600000

=1000000


Contribution margin ratio (CMR):

="\frac{contribution}{Sales}×100"

="\frac{400000}{1000000}×100"

=40%


a) BEP="\frac{FC}{CMR}"

"=\frac{300000}{40\%}=750000"


b)P/V ratio= CMR= 40%


c) Margin of Safety= Actual sales - BEP sales

=1000000-750000

=250000